Given 21 = -5(cos(127) + i sin(127)) 22 = 8(cos(8) + sin(8)) Find the product Z122. Provide your answer below:
Answer #1

Solution:

Using formula,

\small z_{1}z_{2} = r_{1}r_{2}(cos(\theta _{1}+\theta _{2})+isin(\theta _{1}+\theta _{2}))

\small z_{1}z_{2} = -5\times 8(cos(127^{\circ}+8^{\circ})+isin(127^{\circ}+8^{\circ}))

\small z_{1}z_{2} = -40(cos(135^{\circ})+isin(135^{\circ}))

\small z_{1}z_{2} = -40\left ( -\frac{1}{\sqrt{2}}-i\frac{1}{\sqrt{2}} \right )

\small z_{1}z_{2} = \frac{40(1+i)}{\sqrt{2}}

\small z_{1}z_{2} = 20\sqrt{2}(1+i)

Ps: hit the like button